OpenStudy (anonymous):

Assume that you have 3.15 grams of the deadly gas hydrogen cyanide,HCN . What is the volume of the gas at STP?

OpenStudy (anonymous):

The density at STP is ~687 \(\frac{mg}{mL}\) from my appendix. Can you now convert this into a volume based on the mass? :-) density = mass / volume , as you can see from the units above

OpenStudy (anonymous):

Careful though because the density is given in milligrams and you were given 3.15 grams, and you answer should have 3 sig figs.

OpenStudy (anonymous):

I'm not ever sure what you're asking me to do...

OpenStudy (anonymous):

How many grams per milligram (mg)? How many liters per milliliter(mL)?

OpenStudy (anonymous):

3.15

OpenStudy (anonymous):

100

OpenStudy (anonymous):

not3.15

OpenStudy (anonymous):

What? No 3.15g is the mass you have. \(\frac{1}{1000}\)="milli-" prefix

OpenStudy (anonymous):

If you have 687 mg and you convert to grams, how much is this?

OpenStudy (anonymous):

.0687

OpenStudy (anonymous):

Did you mean 0.687? What you wrote was technically incorrect

OpenStudy (anonymous):

.687 is correct.

OpenStudy (anonymous):

\[687 \frac{\cancel{mg}}{\cancel{mL}} * \frac{1 g}{1000 \cancel{mg}} * \frac{1000 \cancel{mL}}{1 L} = ?\] \[\rho = \frac{mass}{volume}\] \[\rho*volume = mass\] \[volume = \frac{mass}{\rho}\] Check: \[L = \huge \frac{g}{(\frac{g}{L})} = \large \frac{g}{1}*\frac{L}{g}\] Make sense now @klag84

OpenStudy (anonymous):

? :-)

OpenStudy (anonymous):

\(\rho\) = rho = international symbol for density that everybody uses, the density is measured relative to a temperature or a phase state (i.e.: solid, liquid, gas) that is dependent on temperature. All I did was Google the STP density of Hydrogen Cyanide and found a table with it on there. :-)

OpenStudy (anonymous):

The only problem was it was in different units than what you're using, so I showed you how to convert above to the units you do need.

OpenStudy (anonymous):

Does this give me the final answer?

OpenStudy (anonymous):

I got the answer

OpenStudy (anonymous):

"What is the volume of the gas at STP?" yes. It depends on the source though, http://www.wolframalpha.com/entities/chemicals/hydrogen_cyanide/bx/aq/hg/ Close but a slightly different density, see?

OpenStudy (anonymous):

Are you sure you weren't given the density of the gas with this problem? If you want to be safe, run both density answers through and give your answer as a range. That's what I'd do if my boss was asking that question to me, give him a range (as low as this to as high as this)